The Magellanic Clouds Newsletter

An electronic exchange on Magellanic Clouds research

Edited by Eva K. Grebel and You-Hua Chu

                                    mcnews@astro.uiuc.edu
                    http://www.astro.uiuc.edu/projects/mcnews/MCNews.html
                            http://www.astro.uni-bonn.de/~mcnews/

No. 49 November 19, 2000

Contents

Abstracts of 12 refereed papers  1
Thesis abstract  9
Abstracts of 3 non-refereed papers 10
Job opportunity  11
Announcement of a school  13

Abstracts of Refereed Papers

A Fractal Analysis of the HI Emission from the Large Magellanic Cloud

Bruce G. Elmegreen^{1}, Sungeun Kim^{2,3}, and Lister Staveley-Smith^{4}

1 IBM Research Division, T.J. Watson Research Center, P.O. Box 218, Yorktown Heights, NY 10598, USA 2 Astronomy Department, University of Illinois at Urbana- Champaign, IL 61801, USA 3 Harvard Smithsonian Center for Astrophysics, 60 Garden Street, Cambridge, MA 02138, USA 4 Australia Telescope National Facility, CSIRO, PO Box 76, Epping, NSW 1710, Australia

Fourier transform power spectra of the distribution of neutral hydrogen emission in the Large Magellanic Cloud is approximately a power law over ∼ 2 decades in length. Power spectra in the azimuthal direction look about the same as the rectilinear spectra. No difference is seen between power spectra of single channel maps and power spectra of either the peak emission map or the integrated emission map at the same location. There is a slight steeping of the average 1D and 2D LMC power spectra at high spatial frequencies. Delta-variance methods also show the same power-law structure.

These results suggest that most of the interstellar medium in the LMC is fractal, presumably the result of pervasive turbulence, self-gravity, and self-similar stirring. The similarity between the channel and integrated maps suggests they cover about the same line-of-sight depth. The slight steepening of the power spectra at high spatial frequency, corresponding to wavelengths smaller than ∼ 100 pc, could mark the transition from large-scale emission that is relatively shallow on the line of sight to small-scale emission that is relatively thick on the line of sight. Such a transition, if real, would provide a method to obtain the thickness of a face-on galactic gas layer.

To check this possibility, three dimensional fractal models are made from the inverse Fourier transform of noise with a power-law cutoff. The models are viewed in projection with a Gaussian density distribution on the line of sight to represent a face-on galaxy disk with finite disk thickness. The density structure from turbulence is simulated in the models by using a log-normal density distribution function with a scale factor dependent on the Mach number. Additional density structure from simulated HI phase transitions is included in some models. After tuning the Mach number, galaxy thickness, and mathematical form of the phase transition, the models can be made to reproduce the observed LMC power spectra, the amplitude of the HI brightness fluctuations, and the probability distribution function for brightness. In all cases, the HI structure arises from a relatively thin layer in the LMC; the thick part of the HI disk has little spatial structure. The large amplitude of the observed intensity variations cannot be achieved by turbulence alone; phase transitions are required.

The character of the fractal HI structure in the LMC is viewed in another way by comparing positive and negative images of the integrated emission. For the isotropic fractal models, these two images have the same general appearance, but for the LMC they differ markedly. The HI is much more filamentary in the LMC than in an isotropic fractal, making the geometric structure of the high-emission regions qualitatively different than the geometric structure of the low-emission (intercloud) regions. The high-emission regions are also more sharply peaked than the low-emission regions, suggesting that compressive events formed the high-emission regions, and expansion events, whether from explosions or turbulence, formed the low-emission regions.

The character of the structure is also investigated as a function of scale using unsharp masks and enlargements with four different resolutions. The circular quality of the low emission regions and the filamentary quality of the high emission regions is preserved on scales ranging from several tens to several hundreds of parsecs.

The spatial scales for sources of turbulent energy input may be illustrated by rms variations in the power spectra with position in the galaxy. This rms decreases from ∼ 0.6 at kpc scales to ∼ 0.25 on ∼ 20 pc scales. The large scale variations are probably from known supershells. The smaller scale variations could be the result of a combination of turbulent cascades from these large-scale energy inputs and additional energy sources with smaller sizes.

Accepted by: The Astrophysical Journal, vol. 548, Feb 10, 2001 For preprints, contact bge@watson.ibm.com Also available from the URL astro-ph/0010578

Supersonic water masers in 30 Doradus

Jacco Th. van Loon^{1} and Albert A. Zijlstra^{2}

1 IoA Cambridge, UK 2 UMIST Manchester, UK

We report on extremely high velocity molecular gas, up to −80 km s^{−1} relative to the ambient medium, in the giant star-formation complex 30 Doradus in the Large Magellanic Cloud (LMC), as observed in new 22 GHz H_{2}O 6_{16} → 5_{23} maser emission spectra obtained with the Mopra radio telescope. The masers may trace the velocities of protostars, and the observed morphology and kinematics indicate that current star formation occurs near the interfaces of colliding stellar-wind blown bubbles. The large space velocities of the protostars and associated gas could result in efficient mixing of the LMC. A similar mechanism in the Milky Way could seed the galactic halo with relatively young stars and gas.

Accepted by: The Astrophysical Journal Letters For preprints, contact jacco@ast.cam.ac.uk

Radio Supernova 1987A at 843 MHz

Lewis Ball^{1}, D. F. Crawford^{1}, R. W. Hunstead^{1}, I. Klamer^{1}, V. J. McIntyre^{1,2}

1 School of Physics, University of Sydney, N.S.W. 2006, Australia 2 Australia Telescope National Facility, CSIRO, PO Box 76, Epping, N.S.W. 2121, Australia

We report the flux densities of the evolving radio source SN1987 A at 843 MHz measured from observations made with the Molonglo Observatory Synthesis Telescope between 1994 September and 2000 May. The radio light curve shows that the rate of increase of the flux density jumped markedly around days 2800-3000 (i.e., in the first half of 1995), and that since then the radio evolution has been remarkably well fitted by a simple linear increase of (62.7±0.5) μJy day^{−1}. We discuss in detail the relationship between the radio light curve and the recent brightening of the system at optical wavelengths.

Accepted by: The Astrophysical Journal

For preprints, go to http://xxx.lanl.gov/abs/astro-ph/0011033
Also available from http://www.physics.usyd.edu.au/∼ball

LSND, SN1987A, and CPT Violation

Hitoshi Murayama^{1,2} and T. Yanagida^{3}

1 Theoretical Physics Group, Ernest Orlando Lawrence Berkeley National Laboratory, University of California, Berkeley, CA 94720, USA 2 Department of Physics, University of California, Berkeley, CA 94720, USA 3 Department of Physics and RESCEU, University of Tokyo, Hongo, Bunkyo-ku, Tokyo 113-0033, Japan

We point out that neutrino events observed at Kamiokande and IMB from SN1987A disfavor the neutrino oscillation parameters preferred by the LSND experiment. For Δm^{2} > 0 (the light side), the electron neutrinos from the neutronization burst would be lost, while the first event at Kamiokande is quite likely to be due to an electron neutrino. For Δm^{2} < 0 (the dark side), the average energy of the dominantly ν̄_{e} events is already lower than the theoretical expectations, which would get aggravated by a complete conversion from ν̄_{μ} to ν̄_{e}. If taken seriously, the LSND data are disfavored independent of the existence of a sterile neutrino. A possible remedy is CPT violation, which allows different mass spectra for neutrinos and anti-neutrinos and hence can accommodate atmospheric, solar and LSND data without a sterile neutrino. If this is the case, Mini-BooNE must run in ν̄ rather than the planned ν mode to test the LSND signal. We speculate on a possible origin of CPT violation.

Submitted to: Physical Review Letters For preprints, contact murayama@lbl.gov

Also available from the URL http://xxx.lanl.gov/abs/hep-ph/0010178

XMM-Newton EPIC Observation of SMC SNR 0102−72.3

M. Sasaki^{1}, T. F. X. Stadlbauer^{1}, F. Haberl^{1}, M. D. Filipović^{1,2,3}, and P. J. Bennie^{4}

1 Max-Planck-Institut für extraterrestrische Physik, Giessenbachstraße, Postfach 1312, 85741 Garching, Germany 2 University of Western Sydney Nepean, P.O. Box 10, Kingswood, NSW 2747, Australia 3 Australia Telescope National Facility, CSIRO, P.O. Box 76, Epping, NSW 2121, Australia 4 University of Leicester, Leicester, LE1 7RH, United Kingdom

Results from observations of the young oxygen-rich supernova remnant SNR 0102−72.3 in the Small Magellanic Cloud during the calibration phase of the XMM-Newton Observatory are presented. Both EPIC-PN and MOS observations show a ringlike structure with a radius of ∼15 arcsec already known from Einstein, ROSAT and Chandra observations. Spectra of the entire SNR as well as parts in the eastern half were analyzed confirming shocked hot plasma in non-uniform ionization stages as the origin of the X-ray emission. The spectra differ in the northeastern and the southeastern part of the X-ray ring, showing emission line features of different strength. The temperature in the northeastern part is significantly higher than in the southeast, reflected by the lines of higher ionization stages and the harder continuum. Comparison to radio data shows the forward shock of the blast wave dominating in the northern part of the SNR, while the southern emission is most likely produced by the recently formed reverse shock in the ejecta. In the case of the overall spectrum of SNR 0102−72.3, the two-temperature non-equilibrium ionization model is more consistent with the data in comparison to the single plane-parallel shock model. The structure of SNR 0102−72.3 is complex due to variations in shock propagation leading to spatially differing X-ray spectra.

Accepted by: Astronomy & Astrophysics Letters, special issue for XMM-Newton For preprints, contact manami@mpe.mpg.de

XMM-Newton Optical Monitor observations of LMC X-3

R.Soria^{1}, K. Wu^{2}, M. J. Page^{1}, and I. Sakelliou^{1}

1 Mullard Space Science Laboratory, University College London, Holmbury St Mary, Dorking, RH5 6NT, UK 2 School of Physics, University of Sydney, NSW 2006, Australia

We study the optical counterpart of the black-hole X-ray binary LMC X−3, by using XMM-Newton/OM observations carried out during a low-hard X-ray state. We derive a better constraint for the temperature, mass and radius of the companion star, and we show that the star is likely to be a ∼ B5 subgiant filling its Roche lobe. Taking into account the effect of X-ray irradiation, we suggest a value f_{M} = (1.5 ± 0.3)M_{⊙} for the mass function in this system, lower than previously thought; we provide a more accurate lower limit to the mass of the compact object.

Accepted by: Astronomy & Astrophysics Letters For preprints, contact rs1@mssl.ucl.ac.uk

Also available from the URL http://www.mssl.ucl.ac.uk/∼rs1/papers/lmcx3-paper1.ps.gz

XMM-Newton EPIC and RGS observations of LMC X-3

K. Wu^{1}, R.Soria^{2}, M. J. Page^{2}, I. Sakelliou^{2}, S. M. Kahn^{3}, and C. P. de Vries^{4}

1 Research Centre for Theoretical Astrophysics, School of Physics, University of Sydney, NSW 2006, Australia 2 Mullard Space Science Laboratory, University College London, Holmbury St Mary, Dorking, RH5 6NT, UK 3 Columbia Astrophysics Laboratory, Columbia University, 550 West 120th Street, New York, NY 10027, USA 4 Space Research Organization of the Netherlands, Sorbonnelaan 2, 3584 CA Utrecht, The Netherlands

We report the results of preliminary analysis of the XMM-Newton EPIC and RGS observations of the candidate black-hole binary LMC X-3 between February and June 2000. The observations covered both the soft and the hard X-ray spectral states. The hard-state spectra were dominated by a powerlaw component with a photon index Γ ≈ 1.9. The soft-state spectra consisted of a thermal component with a multi-color disk temperature T_{in} of ≈ 0.9 keV and a power-law tail with Γ ≈ 2.7. The model in which the X-rays from LMC X-3 in the high-soft state are powered by a strong stellar wind from a massive companion is not supported by the small line-of-sight absorption (n_{H} ∼^{<} 10^{21} cm^{−2}) deduced from the RGS data. The transition from the soft to the hard state appears to be a continuous process associated with the changes in the mass-transfer rate.

Accepted by: Astronomy & Astrophysics Letters For preprints, contact rs1@mssl.ucl.ac.uk

Also available from the URL http://www.mssl.ucl.ac.uk/∼rs1/papers/lmcx3-paper2.ps.gz

Discovery of 101-s Pulsations from AX J0057.4−7325 in the SMC with ASCA

Jun Yokogawa,^{1} Ken’ichi Torii,^{2} Takayoshi Kohmura,^{3} Kensuke Imanishi,^{1} and Katsuji Koyama^{1}

1 Department of Physics, Graduate School of Science, Kyoto University, Sakyo-ku, Kyoto, 606-8502, Japan 2 National Space Development Agency of Japan, 2-1-1 Sengen, Tsukuba, Ibaraki, 305-8505, Japan 3 Department of Earth and Space Science, Graduate School of Science, Osaka University, 1-1 Machikaneyama-cho, Toyonaka, Osaka, 560-0043, Japan

The results from two ASCA observations of AX J0057.4−7325 = RX J0057.3−7325 are presented. Coherent pulsations with a barycentric period of 101.45 ± 0.07 s were discovered in the second observation. The X-ray spectrum was found to be hard (photon index ∼ 0.9) and unchanged through these observations, except for the flux. The ROSAT archival data show that AX J0057.4−7325 exhibits a flux variation with a factor > 10. A discussion on a possible optical counterpart is given.

Accepted by: Publications of the Astronomical Society of Japan

For preprints, contact jun@cr.scphys.kyoto-u.ac.jp
Also available from the URL http://www-cr.scphys.kyoto-u.ac.jp/member/jun/job/

Discovery of the Slowest X-Ray Pulsar in the SMC, AX J0049.5−7323, with ASCA

Jun Yokogawa, Kensuke Imanishi, Masaru Ueno, Katsuji Koyama

Department of Physics, Graduate School of Science, Kyoto University, Sakyo-ku, Kyoto 606-8502, Japan

The discovery of coherent pulsations with ASCA from an X-ray source, AX J0049.5−7323, is reported. The barycentric period was determined to be 755.5 ± 0.6 s, which is the longest among X-ray pulsators in the SMC. The X-ray spectrum has been found to be unchanged through ASCA observations, with a photon index of ∼ 0.8 and a luminosity of ∼ 5 × 10^{35} erg s^{−1} (0.7–10 keV). Archival data of the Einstein and the ROSAT satellites indicate that the flux has been < 10^{−12} erg s^{−1} cm^{−2} (< 5 × 10^{35} erg s^{−1}) for over 20 years and exhibits a variability with a factor of > 10. We argue that AX J0049.5−7323 is an X-ray pulsar with a Be star companion.

Accepted by: Publications of the Astronomical Society of Japan

For preprints, contact jun@cr.scphys.kyoto-u.ac.jp
Also available from the URL http://www-cr.scphys.kyoto-u.ac.jp/member/jun/job/

The Progenitor Masses of Wolf-Rayet Stars

and Luminous Blue Variables Determined from Cluster Turn-offs. II. Results from 12 Galactic Clusters and OB Associations

Philip Massey^{1}, Kathleen DeGioia-Eastwood, and Elizabeth Waterhouse^{2}

1 Lowell Observatory, USA 2 Northern Arizona University, USA

In a previous paper on the Magellanic Clouds we demonstrated that coeval clusters provide a powerful tool for probing the progenitor masses of Wolf-Rayet stars (W-Rs) and Luminous Blue Variables (LBVs). Here we extend this work to the higher metallicity regions of the Milky Way, studying 12 Galactic clusters. We present new spectral types for the unevolved stars, and use these, plus data from the literature, to construct H-R diagrams. We find that all but two of the clusters are highly coeval, with the highest mass stars having formed over a period of less than 1 Myr. The turn-off masses show that at Milky Way metallicities some W-Rs (of early WN type) come from stars with masses as low as 20–25M_{⊙}. Other early-type WNs appears to come from high masses, suggesting a large range of masses evolve through an early-WN stage. On the other hand, WN7 stars are found only in clusters with very high turn-off masses, > 120M_{⊙}. Similarly the LBVs are only found in clusters with the highest turn-off masses, as we found in the Magellanic Clouds, providing very strong evidence that LBVs are a normal stage in the evolution of the most massive stars. The ages of clusters containing WN7s and LBVs can be as young as 1 Myr, and we argue that even so these objects are evolved, and that the young age simply reflects the very high masses that characterize the progenitors of such stars. In particular we show that the archetype LBV η Car appears to be coeval with the rest of the Trumpler 14/16 complex. The three WCs in our sample are all found in clusters with turn-off masses > 70M_{⊙}, whereas in the Magellanic Clouds, the WCs were found in clusters with masses as low as 45M_{⊙}. Whether this difference is significant or due to small-number statistics remains to be seen. The BCs of Galactic W-Rs are hard to establish using the cluster turn-off method, but are consistent with the “standard model” of Hillier.

Accepted by: The Astronomical Journal For preprints, contact massey@lowell.edu

Also available from the URL ftp://ftp.lowell.edu/pub/massey/galwr.ps.gz

Ultraviolet and Optical Observations of OB Associations and Field Stars in the Southwest Region of the Large Magellanic Cloud

Joel Wm. Parker^{1}, Dennis Zaritsky^{2}, Theodore P. Stecher^{3}, Jason Harris^{4}, and Phil Massey^{5}

1 Southwest Research Institute, Suite 426, 1050 Walnut Street, Boulder, Colorado 80302, USA 2 Steward Observatory, Univ. of Arizona, 933 N. Cherry Ave, Tucson, AZ 85721, USA 3 Laboratory for Astronomy and Solar Physics, Code 681, GSFC/NASA, Greenbelt, Maryland 20771, USA 4 UCO/Lick Observatory and Board of Astronomy and Astrophysics, University of California, Santa Cruz, California 95064, USA 5 Lowell Observatory, 1400 W. Mars Hill, Flagstaff, AZ, 86001, USA

Using ultraviolet photometry from the Ultraviolet Imaging Telescope (UIT ) combined with photometry and spectroscopy from three ground-based optical datasets we have analyzed the stellar content of OB associations and field areas in and around the regions N 79, N 81, N 83, and N 94 in the Large Magellanic Cloud. In particular, we compare data for the OB association Lucke-Hodge 2 (LH 2) to determine how strongly the initial mass function (IMF) may depend on different photometric reductions and calibrations. Although the datasets exhibit median photometric differences of up to 30%, the resulting uncorrected IMFs are reasonably similar, typically Γ ∼ −1.6 in the 5–60 M_{⊙} mass range. However, when we correct for the background contribution of field stars, the calculated IMF flattens to Γ = −1.3 ± 0.2 (similar to the Salpeter IMF slope). This change underlines the importance of correcting for field star contamination in determinations of the IMF of star formation regions. It is possible that even in the case of an universal IMF, the variability of the density of background stars could be the dominant factor creating the differences between calculated IMFs for OB associations.

We have also combined the UIT data with the most extensive of these ground-based optical datasets — the Magellanic Cloud Photometric Survey — to study the distribution of the candidate O-type stars in the field. We find a significant fraction, roughly half, of the candidate O-type stars are found in field regions, far from any obvious OB associations [in accord with the suggestions of Garmany, Conti, & Chiosi (1982) for O-type stars in the solar neighborhood]. These stars are greater than 2 arcmin (30 pc) from the boundaries of existing OB associations in the region, which is a distance greater than most O-type stars with typical dispersion velocities will travel in their lifetimes. The origin of these massive field stars (either as runaways, members of low-density star-forming regions, or examples of isolated massive star formation) will have to be determined by further observations and analysis.

Accepted by: The Astronomical Journal (February 2001)

For preprints, contact joel@boulder.swri.edu
Also available from the URL http://www.boulder.swri.edu/∼joel/papers.html

The distance modulus of the Large Magellanic Cloud based on double-mode RR Lyrae stars

G. Kovács^{1}

1Konkoly Observatory, P.O. Box 67, H-1525 Budapest, Hungary

We use double-mode RR Lyrae (RRd) stars from the macho variable star database of the Large Magellanic Cloud (LMC) to estimate its distance, by utilizing photometric data, linear pulsation and stellar atmosphere models. If we set E_{B−V} = 0.11 and [M/H]= −1.5 for LMC, we get M − m = 18.52 mag, which is in agreement with the value obtained from Galactic cluster RRd stars. Although the formal statistical error (standard deviation) of M − m derived from the macho set is only 0.024 mag, the systematic errors are more substantial. These errors are mostly due to the zero point ambiguity of the temperature scale and to the error of the photometric calibration. However, the fact that the distance moduli obtained in this and in our former studies of cluster RRd and Small Magellanic Cloud beat Cepheids agree so well, implies that the only serious source of error is the zero point of the temperature scale, which should not have larger than ±0.10 mag effect on the distance modulus.

Accepted by: Astronomy & Astrophysics Letters For preprints, contact kovacs@konkoly.hu Also available from astro-ph/0011056

Thesis Abstract

A measurement of the abundance of stellar mass dark bodies in the Galactic halo through the search for microlensing phenomena

towards the Magellanic Clouds

T. Lasserre^{1}

1 DAPNIA-SPP, CEA Saclay, 91191 Gif-sur-Yvette, France

Note that the text of the thesis is in French. Here is a translation of the abstract.

Many experimental and theoretical results lead to the conclusion that at least 80 percent of the mass of our Galaxy is dark. Part of this dark matter could be in the form of stellar mass compact objects, the so-called machos; these could be detected using the gravitational microlensing effect. The first generation experiments eros1 and macho have strongly constrained the galactic abundance of objects lighter than 0.01 solar mass to less than 10 percent of the total mass. In parallel, the observation by the macho group of massive candidates (half the Sun’s mass), numerous enough to constitute 50 percent of galactic dark matter, was a further motivation for the eros group to extend this search to stellar mass objects in a second phase, eros2.

The present work deals with the analysis of 25 million stellar light curves in the Large Magellanic Cloud, observed for three years in order to extract the rare microlensing candidates and to measure the galactic halo mass fraction in the form of compact objects. After recalling the motivations of this search and the theoretical context, I describe the eros2 experiment. The observational strategy and the photometric reduction procedures needed to deal with the 1.2 To of data are then presented. A new method to detect microlenses is detailed, as well as a discussion of the still poorly known background light curves. We do not find enough microlensing candidates to explain the galactic rotation curve; this confirms, and improve on previous eros1 and eros2 results. Combining all results from eros allows to exclude that machos with a mass between 10^{−7} and 10 solar mass are important constituents of the galactic halo. This statement agrees with recent results from the macho group, although our interpretations differ, namely on the topics of the location of the lenses, and of a possible contamination of the microlensing sample by background phenomena.

PhD Thesis completed at the Paris-6 University on May 9th, 2000 under the direction of Alain Milsztajn.

For copies or inquiries, contact amilsztajn@cea.fr or lasserre@mpi-hd.mpg.de
Also available from the URL http://www-dapnia.cea.fr/Doc/Publications/Spp/spp00.html
as document DAPNIA/SPP-00-04-T

Abstracts of Non-Refereed Papers

Stellar winds of massive stars

Paul A. Crowther

Dept of Physics & Astronomy, UCL, Gower St, London WC1E 6BT, UK

We review the various techniques through which wind properties of massive stars – O stars, AB supergiants, Luminous Blue Variables (LBVs), Wolf-Rayet (WR) stars and cool supergiants – are derived. The wind momentum-luminosity relation (e.g. Kudritzki et al. 1999) provides a method of predicting mass-loss rates of O stars and blue supergiants which is superior to previous parameterizations. Assuming the theoretical Z^{0.5} metallicity dependence, Magellanic Cloud O star mass-loss rates are typically matched to within a factor of two for various calibrations. Stellar winds from LBVs are typically denser and slower than equivalent B supergiants, with exceptional mass-loss rates during giant eruptions Ṁ = 10^{−3} … 10^{−1}M_{⊙} yr^{−1} (Drissen et al. 2001). Recent mass-loss rates for Galactic WR stars indicate a downward revision of 2–4 relative to previous calibrations due to clumping (e.g. Schmutz 1997), although evidence for a metallicity dependence remains inconclusive (Crowther 2000). Mass-loss properties of luminous (≥ 10^{5}L_{⊙}) yellow and red supergiants from alternative techniques remain highly contradictory. Recent Galactic and LMC results for RSG reveal a large scatter such that typical mass-loss rates lie in the range 10^{−6} … 10^{−4}M_{⊙} yr^{−1}, with a few cases exhibiting ∼ 10^{−3}M_{⊙} yr^{−1}.

To appear in: “The influence of binaries on stellar population studies”, Brussels, Aug 2000 (D. Vanbeveren ed.), Kluwer For preprints, contact pac@star.ucl.ac.uk

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0010581

Local Distance Indicators

Michael Feast^{1}

1 Astronomy Department, University of Cape Town, Rondebosch, 7701, South Africa

A summary is given of the calibration of the Cepheid scale required in the derivation of distances from observations at V and I, as in the HST work on extragalactic Cepheids. There is some evidence that a small metallicity correction may be necessary in deriving distances in this way. The evidence for this comes both from observations of Cepheids of different metallicities and from a comparison of the LMC distance modulus derived from Cepheids and that derived by other methods.

To appear in: IAU Symposium 201, New Cosmological Data and the Values of the Fundamental Parameters, eds. A Lasenby & A Wilkinson

For preprints, contact mwf@artemisia.ast.uct.ac.za
Also available from the URL http://www.arXiv.org/abs/astro-ph/0010590

Not enough stellar mass Machos in the Galactic halo

A. Milsztajn^{1} and T. Lasserre^{1}, for the EROS collaboration

1 DAPNIA-SPP, CEA Saclay, 91191 Gif-sur-Yvette, France

We present an update of results from the search for microlensing towards the Large Magellanic Cloud (lmc) by eros (Expérience de Recherche d’Objets Sombres). We have now monitored 25 million stars over three years. Because of the small number of observed microlensing candidates (four), our results are best presented as upper limits on the amount of dark compact objects in the halo of our Galaxy. We discuss critically the candidates and the possible location of the lenses, halo or lmc. We compare our results to those of the macho group. Finally, we combine these new results with those from our search towards the Small Magellanic Cloud as well as earlier ones from the eros1 phase of our survey. The combined data is sensitive to compact objects in the broad mass range 10^{−7} − 10 M_{⊙}. The derived upper limit on the abundance of stellar mass machos rules out such objects as the dominant component of the Galactic halo if their mass is smaller than 2M_{⊙}.

To appear in: Proceedings of the XIX International Conference on Neutrino Physics and Astrophysics, Sudbury, Canada, June 2000; J. Law, R.W. Ollerhead and J. J. Simpson eds., Nucl. Phys. B (Proc. Suppl.) For preprints, contact AMilsztajn@cea.fr

Also available from the URL http://www-dapnia.cea.fr/Doc/Publications/Spp/spp00.html

as report DAPNIA/SPP-00-26 (a few days from now).

Job Opportunities

POSTDOCTORAL RESEARCH POSITION -

SPACE TELESCOPE SCIENCE INSTITUTE

Applications are invited for a postdoctoral research position at the Space Telescope Science Insti-

tute (http://www.stsci.edu/) to work with Drs. Letizia Stanghellini (www-int.stsci.edu/~lstanghe/),

Richard A. Shaw, and J. Chris Blades on an extensive ground- and HST-based observing program to study extra-Galactic planetary nebulae, stellar evolution, and stellar populations. The successful applicant will be familiar with one or more of the above areas of research, and will also be accomplished in one or more of the following technical areas: hydrodynamic and photo-ionization modeling of gaseous nebulae; the acquisition, calibration, and analysis of ground-based spectroscopic data; and the calibration and analysis of HST imaging and spectroscopy. Familiarity with high-level software languages and astronomical data analysis packages is highly desirable.

Candidates must hold a PhD in astronomy or related field by the date of appointment. The initial appointment is for one year, and is renewable up to 3 years depending on performance and continued funding. There are no restrictions on citizenship. STScI, located on Johns Hopkins University Campus in Baltimore, Maryland, offers an excellent benefit package, competitive salaries, and a stimulating work environment. The minimum salary is $41,900; however, STScI’s pay is commensurate to the year of Ph.D.

Applicants should send a cover letter (mentioning the position applied for), curriculum vitae, list of publications, and a brief (one- or two-page) summary of research interests and accomplishments, and relevant technical expertise to the address below or email to resume@stsci.edu. They should also arrange for three letters of recommendation to be sent to the same address. The position is available immediately, but completed applications received by 15 January 2001 are assured of full consideration. Women and members of minority groups are strongly encouraged to apply. EOE/AA/M/F/D/V.

Space Telescope Science Institute

Attn.: Human Resources

Postdoctoral Research Position - Richard Shaw

3700 San Martin Drive

Baltimore, MD 21218

USA

Questions relating to this position may be directed to Letizia Stanghellini (lstanghe@stsci.edu)

or Dick Shaw (shaw@stsci.edu).
   This information is also posted at http://www.stsci.edu/admin/hres/ads/postdoc12.html.

Meeting Announcement

NAIC – NRAO School on Single-Dish Radio Astronomy – Techniques & Applications

First Announcement

The Arecibo (NAIC) and Green Bank (NRAO) Observatories are organizing the first “NAIC – NRAO School on Single-Dish Radio Astronomy — Techniques & Applications”. The school will take place from June 10 (Sunday) to June 15 (Saturday) 2001 in Arecibo, Puerto Rico. The next school is planned for June 2002 in Green Bank. The school is intended for astronomy graduate students, postdocs, and experts in other astronomical fields who would like to exploit this branch of radio astronomy. It aims at:

• preparing new users of the Arecibo, Green Bank and other single-dish radio telescopes for all aspects of the observing process;

• providing an intensive course on current single-dish capabilities and how to maximize returns from a single-dish telescope;

• providing “hands-on” experience for the participants;

• publishing the proceedings in book form.

The school will consist of an intensive series of lectures (see the preliminary program below), in conjunction with “hands-on” projects. In order to provide as many examples from the “real” singledish world as possible, we invite both lecturers and participants to present posters related to their research experience with single-dishes, emphasizing observational and data-handling techniques.

The number of participants will be limited to approximately 40 people. A registration fee of $125 will include: travel to/from the airport and daily travel to/from the hotels, the welcome reception (held on Sunday, June 10), the social events (Wednesday morning, June 13), the school banquet (Thursday, June 14), and a copy of the proceedings. Financial help should be available for a number of participants.

To find more information about the school and to register, please go to:

                           http://www.naic.edu/meeting.htm

For additional information contact: school@naic.edu.

Preliminary Program:

• Single-Dish Fundamentals:

Introduction: Why single-dishes?

General Theory: Fundamentals of radio astronomy & antennas

Theory of Measurement with Fully-Filled Apertures

The Receiver System

Backends

The Fundamentals of: Continuum, Spectral-Line, Pulsar, Planetary Radar Observations • Single-dish observing branches and associated techniques:

Polarization Continuum Spectral Line Pulsars Planetary Radar Observations Bolometery

• Calibration and Data Reduction:

Calibration Techniques at Radio Wavelengths Calibration Techniques at Millimeter Wavelengths VLBI from the single-dish calibration point of view Reduction & Analysis Techniques

• Special Topics:

Remote Observing The Effects of the Atmosphere and Ionosphere RFI and How to Deal with it (Both in Hardware & Software) Stray Radiation and How to Deal with it Short-spacings correction from the single-dish point of view Focal Plane Arrays The Future for single-dishes

• Hands-On Projects:

AIPS++ “Dish” Lectures & Tutorials Visualization Tools Planning, data acquisition and analysis of simple observations with a real telescope